OpenStudy (anonymous):

How do you solve this problem? An automotive repair center charges $45 for any part of the first hour of labor, and $25 for any part of each additional hour. Which of the following is a correct cost?

OpenStudy (anonymous):

A. C(t) = 145 for 5 < x ≤ 6 B. C(t) = 145 for 6 < x ≤ 7 C. C(t) = 170 for 5 < x ≤ 6 D. C(t) = 170 for 6 < x ≤ 7

OpenStudy (anonymous):

Just see it: In five hours you will have 45 + 4*25 = 45 + 100 = 145. So, if you are between five and six hours, you have to add 25 (any part of each additional hour). So letter A is incorrect. Letter B too, because you will need 45 + 5*25 to 6 hours. It is 170. Letter C is correct. Because you will ned 145 for exactly 5 hours, but when you start the next hours, you will sum 25, and will have 170 until you pass 6 hours. So, the right answer is letter C. Letter D is incorrect because in 7 hours you will have 45 + 6*25 = 195

OpenStudy (texaschic101):

good explaining @zairhenrique ......good job

OpenStudy (anonymous):

thank you for explaining